OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bound by the given curves about the specified line. y=8x^3 y=0 x=1 about x=2 Please Help!

OpenStudy (anonymous):

We have: 

OpenStudy (anonymous):

So then, if we move \(f(x)=8x^3\) over a unit to the left to become g(x) like so: \(g(x)=8(x+1)^3\), then we can simplify the problem to one about revolving the function g(x) about the y-axis!

OpenStudy (dumbcow):

 h = f(x) = 8x^3 r = x+1